• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M Defurne

    Log entry time 16:11:02 on March 06, 2011

    Entry number 350565

    keyword=Day Shift Summary

    SL-TO: Maxime Defurne 3rd: Chunhua Chen

    08:00 Start of Day Shift (setting #12 was at 500mC) 11:00 Stop setting #12 with run 2461. We move spectrometer, change its momentum, change target from LH2 4cm to LH2 15cm. 11:20 Start several runs in order to know if we can use LH2 15cm. And if we can, with how much beam current. 12:00 We decided to run with LH2 15cm under 35uA with a goal 190mC. (First run 2462, the 2464 run crashed 13:30 White alarm appeared on hall A alarm. I made an emtry in logbook. It tells that MF2 Alarm and MF7 Alarm are not connected. We finish run with LH2 15cm. Now we also start calibration run with Al 15cm. 14:00 Switch the target to Al 4cm for calibration run. 14:10 Coda crashed. 2474 JUNK 14:45 After 30-35min (need to reboot several rocs), Coda started working. We started an other calibration run with Aluminium 4cm. 15:45 Start production with LH2 4cm with 80uA.

    setting #12 run 2452 LH2 4cm 80uA 1M Events run 2453 LH2 4cm 80uA 1M Events run 2454 LH2 4cm 80uA 1M Events run 2455 LH2 4cm 80uA 1M Events run 2456 LH2 4cm 80uA 1M Events run 2457 LH2 4cm 80uA 1M Events run 2458 LH2 4cm 80uA 1M Events run 2459 LH2 4cm 80uA 1M Events run 2460 LH2 4cm 80uA 1M Events run 2461 LH2 4cm 80uA 602571 Events

    Total charge for setting #12: 1.365C

    setting #8 run 2462 LH2 15cm 20uA run 2463 LH2 15cm 40uA run 2464 JUNK run 2465 LH2 15cm 35uA 1.1M Events run 2466 LH2 15cm 35uA 1M Events run 2467 LH2 15cm 35uA 1M Events run 2468 LH2 15cm 35uA 1.12M Events run 2469 LH2 15cm 40uA 2M Events run 2470 LH2 15cm 40uA 2M Events run 2471 LH2 15cm 40uA 2M Events run 2472 LH2 15cm 40uA 2M Events Total charge with LH2 4cm: 191mC run 2473 Al 15cm 20uA 590455 Events run 2474 to 2477 JUNK run 2478 Al 4cm 20uA 311039 Events run 2479 BeO 20uA 51372uA

    run 2480 LH2 4cm 80uA in process